OpenStudy (anonymous):

A man travels a distance of 20 miles at 60 miles per hour then returns over the same route at 40 miles per hour. Find his average rate for the round trip in miles. (a) 50 (b)48 (c)47 (d)46

OpenStudy (kropot72):

Total time taken is \[\frac{20}{60}+\frac{20}{40}=\frac{5}{6}\ hour\] Average speed is \[\frac{distance}{time}=\frac{40}{\frac{5}{6}}=\frac{40\times6}{5}\ miles\ per\ hour\]

OpenStudy (anonymous):

240/5 ?

OpenStudy (kropot72):

Correct so far. What is the value of 240/5?

OpenStudy (anonymous):

48?

OpenStudy (kropot72):

That is the correct choice :)

OpenStudy (kropot72):

Note that 50 miles per hour is not the correct answer.

OpenStudy (anonymous):

How come?

OpenStudy (kropot72):

To see why the average speed is not 50 mph see my calculation above.

OpenStudy (anonymous):

Oh...

OpenStudy (kropot72):

The average speed is the total distance travelled (20 + 20 = 40 miles) divided by the time taken (in hours or a fraction of an hour).

OpenStudy (anonymous):

I see. Thank you!

OpenStudy (kropot72):

You're welcome :)